# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 7.2 Encoding equivalent classes
In all this section we consider the integers $1\leq T\leq\ell$, $m\geq 1$ and $1\leq s,a,p\leq 2m\ell$ fixed. To bound $|\mathcal{W}_{\ell,\vec{T}}(s,a,p)|$ we need an efficient way to encode in a unique way an equivalent class $[\gamma]\in\mathcal{W}_{\ell,\vec{T}}(s,a,p)$. To proceed we start by introducing the order of apparition of the elements while we explore chronologically $\gamma$. This will allow us to code $[\gamma]$ by coding what we will call its canonical path. To do so we consider $\boldsymbol{T}:=\{(i,t)\leavevmode\nobreak\ ;\leavevmode\nobreak\ i\in[2m],\leavevmode\nobreak\ t\in[T_{i}]\}$ the set of times, that we consider ordered canonically, i.e. for $i<i^{\prime}$ we have for any $t\in[T_{i}]$ and $t^{\prime}\in[T_{i^{\prime}}]$ $(i,t)\prec(i^{\prime},t^{\prime})$ and for $1\leq t<t^{\prime}\leq T_{i}$ we have $(i,t)\prec(i,t^{\prime})$. For $(i,t)\neq(1,1)$ we denote by $(i,t)^{-}$ the greatest element of $\boldsymbol{T}$ strictly smaller than $(i,t)$. For a time $(i,t)\in\boldsymbol{T}$ we set $\gamma_{i,t}:=(x_{i,t},{y}_{i,t},{x}_{i,t+1})$. For all $x\in X_{\gamma}$ (resp. $y\in Y_{\gamma}$) we denote by $\bar{x}$ (resp. $\bar{y}$) the order of apparition of $x$ (resp. of $y$) in $\gamma$. We consider $\bar{\gamma}:=(\bar{x}_{1,1},\bar{y}_{1,1},\ell_{1,1},\bar{x}_{1,2},...,\bar{y}_{i,t},\ell_{i,t},\bar{x}_{i,t+1},...,\bar{y}_{2m,T_{2m}},\ell_{2m,T_{2m}},\bar{x}_{2m,T_{2m}+1})$ the canonical path and knowing $\bar{\gamma}$ is equivalent to know $[\gamma]$. In fact, we do not have necessarily $\bar{\gamma}\in[\gamma]$, for instance $\bar{\gamma}_{i}=(\bar{x}_{i,1},\bar{y}_{i,1},\ell_{i,1},\bar{x}_{i,2},...)$ can be tangled. However, if $K_{\bar{\gamma}}$ is constructed as we did for $K_{\gamma}$ then for all $\gamma^{\prime}\in[\gamma]$, the graph $K_{\gamma^{\prime}}$ is isomorphic to $K_{\bar{\gamma}}$.
@@
$\bar{c}(\bar{x}):=\min\{\bar{x}\leavevmode\nobreak\ ;\leavevmode\nobreak\ x\in\operatorname{cc}(x^{\prime})\}.$
@@
Notice that at a first time that is not a merging time, the $x\in X_{\gamma}$ discovered gives the color of its connected component. If $(i,t)$ is an important time, we first mark it with the vector $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is the next time we step outside of the tree $\Gamma_{i,t}=\Gamma_{(i,t)^{-}}$. If for the remaining time of the era $i$ the path $\gamma$ stays in $\Gamma_{i,t}$ then we set $\tau=T_{i}+1$. If we have $\tau\neq T_{i}+1$ then $(i,\tau)$ is the next important, merging or first time. If $(i,t)$ is a merging time we also mark it $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is once again the next important, merging or first time. For these two marked times, in the time in between $(i,t)$ and $(i,\tau)$ we only go through seen vertices of $K_{\gamma}$ and seen vertices of $\Gamma_{i,t}$ and therefore one can recover $\{\bar{\gamma}_{i,s}:\leavevmode\nobreak\ t\leq s\leq\tau\}$. Finally for $i\in[2m]$ odd, we call first outgoing time $(i,t)$ the first time occurring that is either a first or important time and mark it with $(\vec{x}_{1,1},\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is the next important time. If a first outgoing time is also a first time, in the time in between $(i,t)$ and $(i,\tau)$ we only see first times. If for $t\leq s\leq T_{i}$ we have only first times we set $\tau=T_{i}+1$. Therefore knowing the first outgoing, important and merging times and their marks we can recover all $\bar{\gamma}$.
@@
We can describe the recovering of $\bar{\gamma}$ as follows. One has to segment the recovering with each eras $i$. If the time $(1,1)$ is not marked then it a first time, therefore we can set that $\bar{x}_{1,1}=1$, $\bar{c}(1)=\bar{c}(\bar{x}_{1,1})=1$, $\bar{y}_{1,1}=1$, $\bar{x}_{1,2}=2$ and finally $\bar{c}(2)=\bar{c}(\bar{x}_{1,2})=2$. Therefore we know that $\bar{\gamma}_{1,1}=((1,1),1,(2,2))$. If now $(1,1)$ is an important time, we still have $\bar{x}_{1,1}=1$, $\bar{c}(1)=1$ and $\bar{y}_{1,1}=1$ but then we have $\bar{c}(2)=1$ and we know the value of $\bar{x}_{1,2}$ from the mark on $(1,1)$. Since $\Gamma_{1,1}=\Gamma_{1,0}$ has no edge, necessarily the next time we step outside the tree is the next time, i.e the time $(1,2)$ is the next first or important time. We start by explaining, for an era $j=2i$ how to recover $\bar{\gamma}_{j,s}$ for $1\leq s\leq t$ where $(j,t)$ is the first marked time but not first outgoing time of the era $j$ considering that $\bar{\gamma}_{j^{\prime},s^{\prime}}$ for $(1,1)\prec(j^{\prime},s^{\prime})\prec(j,1)$ are known. Let $(2i,t)$ be the first marked time and let $(\bar{y}_{2i,t},\vec{x}_{2i,t+1},\vec{x}_{2i,\tau})$ be its mark. If $\bar{y}_{2i,t}$ (or $\bar{c}(\bar{x}_{2i,t+1})$) has never been seen during the preceding eras then it means $(2i,t)$ is an important time and in the time in between $(2i,1)$ and $(2i,t)^{-}$ we only have first times. We can recover then $\bar{\gamma}_{2i,s}$ for $1\leq s\leq t$. If now $\bar{y}_{2i,t}$ or $\bar{c}(\bar{x}_{2i,t+1})$ has been seen before, then $(2i,t)$ is the merging time and once again we know that we only have first times between $(2i,1)$ and $(2i,t)^{-}$. Now we consider an era $j=2i+1$, $(2i+1,t)$ its first outgoing time and $(\vec{x}_{2i+1,1},\bar{y}_{2i+1,t},\vec{x}_{2i+1,t},\vec{x}_{2i+1,\tau})$ its mark. Then there exists only one path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,\tau})$ and even though $\tau$ is not indexed in the mark it is given by the length of this path. Now if there exists another marked time $(2i+1,t^{\prime})$ in the same era $12i+$ 1, it is necessarily an important time. If we assume $\tau\neq t^{\prime}$, we have $\bar{\gamma}_{2i+1,s}$ for $1\leq s\leq\tau$ given by the only path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,\tau})$. We then recover $\bar{\gamma}_{2i+1,s}$ for $\tau+1\leq s\leq t-1$ from the fact that $\{(2i+1,s):\leavevmode\nobreak\ \tau+1\leq s\leq t^{\prime}-1\}$ are first times. If $\tau=t^{\prime}$ then $\bar{\gamma}_{2i+1,s}$ for $1\leq s\leq\tau=t^{\prime}$ is given by the only path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,t^{\prime}})$. If there is no other marked time in the era $2i$ then we only have first times (and $\tau=T_{2i+1}+1$). We now consider $(i,t)\prec(i,t^{\prime})$ two consecutive marked times of a same era $i$ and we suppose $(\bar{\gamma}_{j,s})_{(j,s)\prec(i,t)}$ is known. Let $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ be the mark on $(i,t)$ and $(\bar{y}_{i,t^{\prime}},\vec{x}_{i,t^{\prime}+1},\vec{x}_{i,\tau^{\prime}})$ the one on $(i,t^{\prime})$. We have only one path (possibly empty) possible between $\operatorname{cc}(x_{i,t+1})$ and $\operatorname{cc}(x_{i,\tau})$. The length of this path and the path itself gives us the value of $\tau$ and $\bar{\gamma}_{i,s}$ for $t\leq s\leq\tau$. If $\tau=t^{\prime}$ then we have recovered $\bar{\gamma}_{i,s}$ for $t\leq s\leq t^{\prime}$. Otherwise all $(i,s)$ for $\tau\leq s\leq t^{\prime}$ are necessarily first times.
@@
To proceed we consider each era $i$ separately. We use the fact that $\gamma_{i}$ being tangle-free, if we explore a cycle during the era $i$ we know it to be the only one. For $i\in[2m]$ let $(i,t_{0})$ be the smallest time such that $\operatorname{cc}(x_{i,t_{0}+1})\in\{\operatorname{cc}(x_{i,1}),\operatorname{cc}(x_{i,2}),...,\operatorname{cc}(x_{i,t_{0}})\}$, if it exists. Let $1\leq\sigma\leq t_{0}$ be such that $\operatorname{cc}(x_{i,t_{0}+1})=\operatorname{cc}(x_{i,\sigma})$. We denote by $\mathrm{C}_{i}:=(\bar{\operatorname{cc}}(x_{i,\sigma}),\bar{\operatorname{cc}}(x_{i,\sigma+1}),...,\bar{\operatorname{cc}}(x_{i,t_{0}+1}))$ and $\mathrm{C}_{i}$ is the only cycle visited by $\gamma_{i}$. We define the short cycling time as the greatest important time $(i,t)$ smaller then $(i,t_{0})$ in $(\boldsymbol{T},\prec)$ and $(i,\hat{t})$ the smallest time $(i,\hat{t})\succeq(i,\sigma)$ verifying that $\operatorname{cc}(x_{i,\hat{t}+1})\notin\mathrm{C}_{i}$. As convention we take $\hat{t}=T_{i}+1$ if $\gamma_{i}$ stays in $\mathrm{C}_{i}$ for the remaining time of the era $i$. If $\hat{t}\geq t_{0}+2$ then the cycle $\mathrm{C}_{i}$ is visited more then once, even though it may not finish the cycle after the first visit. We replace the mark on $(i,t)$ with $(\bar{y}_{i,t},\vec{x}_{i,t+1},\sigma,\hat{t},\vec{x}_{i,\tau})$ where now $(i,\tau)$ is the next important time after $(i,\hat{t})$. We take as convention that $\tau=T_{i}+1$ if $\gamma_{i}$ stays in $\Gamma_{(i,t)}$ for the remaining time of the era $i$. During the era $i$ we have only one short cycling time. The important times $(i,t^{\prime})$ that verify $1\leq t^{\prime}\leq t$ or $\tau\leq t^{\prime}\leq T_{i}$ are called long cycling time. The important times in the time in between $(i,t)$ and $(i,\tau)$ are called superfluous and are not marked anymore. We can now prove 7.2.
Proof of 7.2.
@@
7.3 Contributions of the matrix $Q$ for a fixed path and proof of 7.1
In this section we fix $[\gamma]\in\mathcal{W}_{\ell,\vec{T}}(s,a,p)$ and we denote by $a_{1}$ the number of inconsistent edges $e\in\mathrm{A}_{\gamma}$ of multiplicity $1$ (see 6.5).
@@
Let $m\geq 1$ and $1\leq\ell\leq\frac{\ln(N)}{6\ln(d)}$ be fixed integers. There exists $c,c(n)\geq 1$ such that
$\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\leq c(n)N^{s-p}(cm\ell)^{a_{1}+3(m+g)+p}\rho_{n,\#}^{2m(\ell+1)}.$
@@
Let $\mathcal{G}=(\mathcal{X}\sqcup\mathcal{Y},\mathcal{E})$ be a bi-partite connected and directed graph in the sense of 7.6 and $\mathcal{T}$ an underlying tree. If there is a double edge in $\mathcal{G}$ we denote it $\𝒯(e,e^{\prime})\in\mathcal{T}\times(\mathcal{E}\backslash\mathcal{T})$. Let $(\ell(f))_{f\in\mathcal{E}}$ be a family of integers. We define
$Z(\mathcal{G}):=\sum_{\begin{subarray}{c}y_{u}\in[N]\\ u\in\mathcal{\mathcal{Y}}\end{subarray}}\sum_{\begin{subarray}{c}x_{v}\in[N]\\ v\in\mathcal{\mathcal{X}}\end{subarray}}\prod_{\begin{subarray}{c}f\in\mathcal{E}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{y_{u}x_{v}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ L_{\mathcal{T}}:=\sum_{f\in\mathcal{T}\backslash\{e\}}\ell(f),$
@@
$Z(\mathcal{G})\leq N\operatorname{Tr}_{N}\left((Q^{\ell(e)}{Q^{\ast}}^{\ell(e^{\prime})}\right)\delta^{L_{\mathcal{T}}}.$
@@
We then bound the sum over $y_{u_{0}}\in[N]$ by $\delta^{\ell(e_{0})}$ and apply the induction to $\mathcal{G}^{\prime}=(\mathcal{X}\sqcup(\mathcal{Y}\backslash\{u_{0}\}),\mathcal{E}\backslash\{e_{0}\})$. In the case where we have $\mathrm{d}^{\mathrm{o}}_{\mathcal{T}}(v_{0})=1$ we similarly obtain
@@
We then bound the sum over $x_{v_{0}}\in[N]$ by $\delta^{\ell(e_{0})}$ and apply the induction to $\mathcal{G}^{\prime}=((\mathcal{X}\backslash\{v_{0}\})\sqcup\mathcal{Y},\mathcal{E}\backslash\{e_{0}\})$. ∎
@@
$q(\gamma)=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t},x_{i,t+1}}=\prod_{\operatorname{cc}\in\chi}\prod_{\begin{subarray}{c}f\in\mathcal{E}_{\operatorname{cc}}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{x_{u},x_{v}}$
where $\chi:=X_{\gamma}/\sim$ are the connected components of $K_{\gamma}$ and for $\operatorname{cc}\in\chi$ we considered the graph $\mathcal{G}_{\operatorname{cc}}:=(\mathcal{V}_{\operatorname{cc}},\mathcal{E}_{\operatorname{cc}})$ defined by $\mathcal{V}_{\operatorname{cc}}:=\{\bar{x};\leavevmode\nobreak\ x\in\operatorname{cc}\}\sqcup\{\bar{y};\leavevmode\nobreak\ \exists(i,t)\in\boldsymbol{T},\leavevmode\nobreak\ x\in\operatorname{cc};\leavevmode\nobreak\ (y_{i,t},x_{i,t+1})=(y,x)\}=:\mathcal{X}_{\operatorname{cc}}\sqcup\mathcal{Y}_{\operatorname{cc}}$ and $\mathcal{E}_{\operatorname{cc}}$ is the multi-set defined by $\mathcal{E}_{\operatorname{cc}}:=\{(u,v)\in\mathcal{V}_{\operatorname{cc}}^{2};\leavevmode\nobreak\ \leavevmode\nobreak\ \exists(i,t)\in\boldsymbol{T};\leavevmode\nobreak\ \leavevmode\nobreak\ (\bar{y}_{i,t},\bar{x}_{i,t+1})=(u,v)\}$. The notations are the one introduced in the proof of 7.2 and the graphs $(\mathcal{G}_{\operatorname{cc}})_{\operatorname{cc}\in\chi}$ are directed, weakly connected and bi-partite in the sense of 7.6. In the writing of $q(\gamma)$ above we have considered that for $u=\bar{x}$ we set $x_{u}=x$. We consider $\mathcal{T}_{\operatorname{cc}}$ the covering tree obtained by constructing the tree while we explore $\gamma$ chronologically and we set $\mathcal{R}_{\operatorname{cc}}:=\mathcal{G}_{\operatorname{cc}}\backslash\mathcal{T}_{\operatorname{cc}}$. We introduce $\chi_{\mathrm{d}}\subset\chi$ the set of connected component $\operatorname{cc}\in\chi$ such that there exists $(e(\mathrm{cc}),e^{\prime}(\mathrm{cc}))$ a double edge in $\mathcal{G}_{\operatorname{cc}}$. We consider the graphs
$\mathcal{G}=(\mathcal{X}\sqcup\mathcal{Y},\mathcal{E}):=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{G}_{\operatorname{cc}},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{F}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{T}_{\operatorname{cc}}\backslash\{e(\operatorname{cc})\}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{R}:=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathcal{D}:=\bigsqcup_{\operatorname{cc}\in\chi}\{e(\operatorname{cc}),e^{\prime}(\operatorname{cc})\}$
@@
$F:=|\mathcal{F}|\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ R:=|\mathcal{R}|.$ (31)
In order to prove 7.9, we provide an upper bound for $F$.
@@
We use the previous lemmas to prove 7.5.
Proof of 7.5.
For a finite set $\mathcal{S}$ and an integer $I$ we introduce
$\mathcal{L}^{\prime}(\mathcal{S},I):=\left\{(\ell(s))_{s\in\mathcal{S}}\in[\ell]^{\mathcal{S}}:\leavevmode\nobreak\ \sum_{s\in\mathcal{S}}\ell(s)=I\right\}.$
@@
We give the detail of the upper bound for $q$ defined by (29) and we precise along the proof where the reasoning needs to be adapted for $q_{r}$. We bound our sum as follows
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $=\sum_{\gamma^{\prime}\sim\gamma}\prod_{\operatorname{cc}\in\chi}\prod_{\begin{subarray}{c}f\in\mathcal{E}_{\operatorname{cc}}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{x^{\prime}_{u},x^{\prime}_{v}}\leq\sum_{\begin{subarray}{c}y_{u},x_{v}\in[N]\\ (u,v)\in\mathcal{Y}\times\mathcal{X}\end{subarray}}\sum_{(\ell(f))_{f\in\mathcal{E}}\in\mathcal{L}^{\prime}(\mathcal{E},M)}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{E}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$ (32)
where we set $M=2m(\ell+1)-T$ for
$T:=\sum_{i=1}^{2m}T_{i}.$
For a fixed choice of $\{(y_{u},x_{v}):\leavevmode\nobreak\ (u,v)\in\mathcal{Y}\times\mathcal{X}\}$, we rewrite and bound the second sum above as follows
$\sum_{\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{E}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$ $=\sum_{L=0}^{M}\sum_{\begin{subarray}{c}\ast\ast\end{subarray}}\prod_{\operatorname{cc}\in\chi}\prod_{f\notin(\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\})}(Q^{\ell(f)})_{y_{u}x_{v}}\sum_{\ast\ast\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in(\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\})}(Q^{\ell(f)})_{y_{u}x_{v}}$
[…]
$\sum_{\ast\ast\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in(\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\})}(Q^{\ell(f)})_{y_{u}x_{v}}\leq\binom{M-L+R-1}{R-1}\kappa^{M-L+R}\leq(2\kappa)^{M-L+R-1}\kappa.$ (33)
Implementing the previous bound in (32)

### v4

@@ 8.2 Encoding equivalent classes
[…]
@@
$\bar{c}(\bar{x}):=\min\{\bar{x}\penalty 10000\ ;\penalty 10000\ x\in\operatorname{cc}(x^{\prime})\}.$
@@
Notice that at a first time that is not a merging time, the $x\in X_{\gamma}$ discovered gives the color of its connected component. If $(i,t)$ is an important time, we first mark it with the vector $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is the next time we step outside of the tree $\Gamma_{i,t}=\Gamma_{(i,t)^{-}}$. If for the remaining time of the era $i$ the path $\gamma$ stays in $\Gamma_{i,t}$ then we set $\tau=T_{i}+1$. If we have $\tau\neq T_{i}+1$ then $(i,\tau)$ is the next important, merging or first time. If $(i,t)$ is a merging time we also mark it $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is once again the next important, merging or first time. For these two marked times, in the time in between $(i,t)$ and $(i,\tau)$ we only go through seen vertices of $K_{\gamma}$ and seen vertices of $\Gamma_{i,t}$ and therefore one can recover $\{\bar{\gamma}_{i,s}:\penalty 10000\ t\leq s\leq\tau\}$. Finally for $i\in[2m]$ odd, we call first outgoing time $(i,t)$ the first time occurring that is either a first or important time and mark it with $(\vec{x}_{1,1},\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is the next important time. If a first outgoing time is also a first time, in the time in between $(i,t)$ and $(i,\tau)$ we only see first times. If for $t\leq s\leq T_{i}$ we have only first times we set $\tau=T_{i}+1$. Therefore knowing the first outgoing, important and merging times and their marks we can recover all $\bar{\gamma}$.
@@
We can describe the recovering of $\bar{\gamma}$ as follows. One has to segment the recovering with each eras $i$. If the time $(1,1)$ is not marked then it a first time, therefore we can set that $\bar{x}_{1,1}=1$, $\bar{c}(1)=\bar{c}(\bar{x}_{1,1})=1$, $\bar{y}_{1,1}=1$, $\bar{x}_{1,2}=2$ and finally $\bar{c}(2)=\bar{c}(\bar{x}_{1,2})=2$. Therefore we know that $\bar{\gamma}_{1,1}=((1,1),1,(2,2))$. If now $(1,1)$ is an important time, we still have $\bar{x}_{1,1}=1$, $\bar{c}(1)=1$ and $\bar{y}_{1,1}=1$ but then we have $\bar{c}(2)=1$ and we know the value of $\bar{x}_{1,2}$ from the mark on $(1,1)$. Since $\Gamma_{1,1}=\Gamma_{1,0}$ has no edge, necessarily the next time we step outside the tree is the next time, i.e the time $(1,2)$ is the next first or important time. We start by explaining, for an era $j=2i$ how to recover $\bar{\gamma}_{j,s}$ for $1\leq s\leq t$ where $(j,t)$ is the first marked time but not first outgoing time of the era $j$ considering that $\bar{\gamma}_{j^{\prime},s^{\prime}}$ for $(1,1)\prec(j^{\prime},s^{\prime})\prec(j,1)$ are known. Let $(2i,t)$ be the first marked time and let $(\bar{y}_{2i,t},\vec{x}_{2i,t+1},\vec{x}_{2i,\tau})$ be its mark. If $\bar{y}_{2i,t}$ (or $\bar{c}(\bar{x}_{2i,t+1})$) has never been seen during the preceding eras then it means $(2i,t)$ is an important time and in the time in between $(2i,1)$ and $(2i,t)^{-}$ we only have first times. We can recover then $\bar{\gamma}_{2i,s}$ for $1\leq s\leq t$. If now $\bar{y}_{2i,t}$ or $\bar{c}(\bar{x}_{2i,t+1})$ has been seen before, then $(2i,t)$ is the merging time and once again we know that we only have first times between $(2i,1)$ and $(2i,t)^{-}$. Now we consider an era $j=2i+1$, $(2i+1,t)$ its first outgoing time and $(\vec{x}_{2i+1,1},\bar{y}_{2i+1,t},\vec{x}_{2i+1,t},\vec{x}_{2i+1,\tau})$ its mark. Then there exists only one path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,\tau})$ and even though $\tau$ is not indexed in the mark it is given by the length of this path. Now if there exists another marked time $(2i+1,t^{\prime})$ in the same era $2i+1$, it is necessarily an important time. If we assume $\tau\neq t^{\prime}$, we have $\bar{\gamma}_{2i+1,s}$ for $1\leq s\leq\tau$ given by the only path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,\tau})$. We then recover $\bar{\gamma}_{2i+1,s}$ for $\tau+1\leq s\leq t-1$ from the fact that $\{(2i+1,s):\penalty 10000\ \tau+1\leq s\leq t^{\prime}-1\}$ are first times. If $\tau=t^{\prime}$ then $\bar{\gamma}_{2i+1,s}$ for $1\leq s\leq\tau=t^{\prime}$ is given by the only path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,t^{\prime}})$. If there is no other marked time in the era $2i$ then we only have first times (and $\tau=T_{2i+1}+1$). We now consider $(i,t)\prec(i,t^{\prime})$ two consecutive marked times of a same era $i$ and we suppose $(\bar{\gamma}_{j,s})_{(j,s)\prec(i,t)}$ is known. Let $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ be the mark on $(i,t)$ and $(\bar{y}_{i,t^{\prime}},\vec{x}_{i,t^{\prime}+1},\vec{x}_{i,\tau^{\prime}})$ the one on $(i,t^{\prime})$. We have only one path (possibly empty) possible between $\operatorname{cc}(x_{i,t+1})$ and $\operatorname{cc}(x_{i,\tau})$. The length of this path and the path itself gives us the value of $\tau$ and $\bar{\gamma}_{i,s}$ for $t\leq s\leq\tau$. If $\tau=t^{\prime}$ then we have recovered $\bar{\gamma}_{i,s}$ for $t\leq s\leq t^{\prime}$. Otherwise all $(i,s)$ for $\tau\leq s\leq t^{\prime}$ are necessarily first times.
@@
[…]
Proof of ˜8.1.
@@
8.3 Contributions of the matrix $Q$ for a fixed path and proof of Section˜8.1
In this section we fix the integers $\ell,m\geq 1$ such that
$1\leq\ell\leq\frac{\ln(N)}{6(\ln(d)\vee 2\ln(\delta\vee 1))}\text{ and }1\leq m\leq\ln(N).$ (34)
We also fix the integers $0\leq s,a,p\leq 2m\ell$ and consider $\gamma]\in\mathcal{W}_{\ell,\vec{T}}(s,a,p)$ for some $\vec{T}\in[\ell]^{2m}$ and we denote by $a_{1}$ the number of inconsistent edges $e\in\mathrm{A}_{\gamma}$ of multiplicity $1$ (see Section˜7.3).
@@
There exists $c,c(n)\geq 1$ such that
$\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{x}}(\gamma^{\prime})\leq c(n)N^{s-p}(cm\ell)^{a_{1}+3(m+g)+p}\rho_{n,\mathrm{x}}^{2m\ell}.$
@@
Let $\mathcal{G}=(\mathcal{X}\sqcup\mathcal{Y},\mathcal{E})$ be a bi-partite connected and directed graph in the sense of ˜8.1 and $\mathcal{T}$ an underlying tree. If there is a double edge in $\mathcal{G}$ we denote it $(e,e^{\prime})\in\mathcal{T}\times(\mathcal{E}\backslash\mathcal{T})$. Let $(\ell(f))_{f\in\mathcal{E}}$ be a family of integers. We define
$Z(\mathcal{G}):=\sum_{\begin{subarray}{c}y_{u}\in[N]\\ u\in\mathcal{\mathcal{Y}}\end{subarray}}\sum_{\begin{subarray}{c}x_{v}\in[N]\\ v\in\mathcal{\mathcal{X}}\end{subarray}}\prod_{\begin{subarray}{c}f\in\mathcal{E}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{y_{u}x_{v}}\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ L_{\mathcal{T}}:=\sum_{f\in\mathcal{T}\backslash\{e\}}\ell(f),$
@@
$Z(\mathcal{G})\leq N\operatorname{Tr}_{N}\left(Q^{\ell(e)}{Q^{\ast}}^{\ell(e^{\prime})}\right)\delta^{L_{\mathcal{T}}}.$
@@
[…]
@@
[…]
@@
$q_{\mathrm{s}}(\gamma)=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t},x_{i,t+1}}=\prod_{\operatorname{cc}\in\chi}\prod_{\begin{subarray}{c}f\in\mathcal{E}_{\operatorname{cc}}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{x_{u},x_{v}}$
where $\chi:=X_{\gamma}/\sim$ are the connected components of $K_{\gamma}$ and for $\operatorname{cc}\in\chi$ we considered the graph $\mathcal{G}_{\operatorname{cc}}:=(\mathcal{V}_{\operatorname{cc}},\mathcal{E}_{\operatorname{cc}})$ defined by $\mathcal{V}_{\operatorname{cc}}:=\{\bar{x};\penalty 10000\ x\in\operatorname{cc}\}\sqcup\{\bar{y};\penalty 10000\ \exists(i,t)\in\boldsymbol{T},\penalty 10000\ x\in\operatorname{cc};\penalty 10000\ (y_{i,t},x_{i,t+1})=(y,x)\}=:\mathcal{X}_{\operatorname{cc}}\sqcup\mathcal{Y}_{\operatorname{cc}}$ and $\mathcal{E}_{\operatorname{cc}}$ is the multi-set defined by $\mathcal{E}_{\operatorname{cc}}:=\{(u,v)\in\mathcal{V}_{\operatorname{cc}}^{2};\penalty 10000\ \penalty 10000\ \exists(i,t)\in\boldsymbol{T};\penalty 10000\ \penalty 10000\ (\bar{y}_{i,t},\bar{x}_{i,t+1})=(u,v)\}$. The notations are the one introduced in the proof of ˜8.1 and the graphs $(\mathcal{G}_{\operatorname{cc}})_{\operatorname{cc}\in\chi}$ are directed, weakly connected and bi-partite in the sense of ˜8.1. In the writing of $q(\gamma)$ above we have considered that for $u=\bar{x}$ we set $x_{u}=x$. We consider $\mathcal{T}_{\operatorname{cc}}$ the covering tree obtained by constructing the tree while we explore $\gamma$ chronologically and we set $\mathcal{R}_{\operatorname{cc}}:=\mathcal{G}_{\operatorname{cc}}\backslash\mathcal{T}_{\operatorname{cc}}$. We introduce $\chi_{\mathrm{d}}\subset\chi$ the set of connected component $\operatorname{cc}\in\chi$ such that there exists $(e(\mathrm{cc}),e^{\prime}(\mathrm{cc}))$ a double edge in $\mathcal{G}_{\operatorname{cc}}$. We consider the graphs
[…]
@@
$F:=|\mathcal{F}|\penalty 10000\ \text{and}\penalty 10000\ R:=|\mathcal{R}|.$ (35)
In order to prove ˜8.7, we provide an upper bound for $F$.
@@
We use the previous lemmas to prove ˜8.4.
Proof of ˜8.4.
@@
We give the detail of the upper bound for $q_{\mathrm{s}}$ defined by (32) and we precise along the proof where the reasoning needs to be adapted for $q_{\mathrm{c}}$.
For $\gamma^{\prime}=(x^{\prime}_{1,1},y^{\prime}_{1,1},\ell_{1,1}^{\prime},x^{\prime}_{1,2},...,x^{\prime}_{2m,T_{2m}+1})\sim\gamma$ and considering the weighted graph defined above we set
$L_{\mathcal{R}}(\gamma^{\prime})=\sum_{f\in\mathcal{R}}\ell^{\prime}(f).$
We first bound our sum as follows